# Sequential Monte-Carlo testing by betting

### v4

@@ 4.2 A theoretically grounded choice for binomial parameter $p$
Comparing the binomial strategy with the extreme ones in Section 2, the parameter $p$ can be interpreted as the aggressiveness with which we bet. The lower the $p$, the more aggressive we bet on the alternative. In the following, we introduce a choice of $p$ that adapts to the level $\alpha\in(0,1)$ the considered hypothesis is tested at and which offers a good compromise between aggressiveness and safety. First, we need the following lemma.
@@
Let $p=1/N$ for some $N\in\mathbb{N}$, and let $k_{N}$ denote the realized number of losses after $N$ permutations. If $p_{O}\leq p$ for some $B\geq N$ and $k_{N}\geq 1$, then there exists $\tilde{B}\leq B$ such that $p=k_{\tilde{B}}/\tilde{B}$.
Proof.
Let $\tilde{B}=\min\{b\in\mathbb{N}:k_{b}/b\leq p,k_{b}\geq 1\}$. Note that $k_{\tilde{B}}/(Nk_{\tilde{B}})=p$. Therefore, $Nk_{\tilde{B}}\leq\tilde{B}$ (otherwise $k_{\tilde{B}}/\tilde{B}>p$) and thus $k_{Nk_{\tilde{B}}}\leq k_{\tilde{B}}$. Since $k_{Nk_{\tilde{B}}}/(Nk_{\tilde{B}})\leq k_{\tilde{B}}/(Nk_{\tilde{B}})=p$ and $k_{Nk_{\tilde{B}}}\geq k_{N}\geq 1$, we have $\tilde{B}=Nk_{\tilde{B}}=k_{\tilde{B}}/p$. ∎
The following proposition applies Lemma 4.3 to find a $p$ such that the hypothesis is always rejected if the classical permutation p-value $\mathfrak{p}^{\mathrm{perm}}_{B}$ is less than or equal to $p$ for some $B\in\mathbb{N}$.
Proposition 4.4.
Let $N=\lceil{\sqrt{2\pi e^{1/6}}/\alpha}\rceil$ and $p=1/N$. If $\mathfrak{p}^{\mathrm{perm}}_{B}\leq p$ for some $B\in\mathbb{N}$, then there exists $\tilde{B}\leq B$ such that $W_{\tilde{B}}^{p}(k_{\tilde{B}})\geq 1/\alpha$.
Proof.
First, assume that $k_{N}\geq 1$. Since $k_{B}/B\leq\mathfrak{p}^{\mathrm{perm}}_{B}\leq p$ for all $B\in\mathbb{N}$, Lemma 4.3 shows that there exists $\tilde{B}\in\mathbb{N}$ such that $p=k_{\tilde{B}}/\tilde{B}=1/N$. Using the first bound of Proposition 4.2, we obtain
$W_{\tilde{B}}^{p}(k_{\tilde{B}})$ $\geq\tilde{B}\sqrt{(\tilde{B}+1)/(k_{\tilde{B}}(\tilde{B}-k_{\tilde{B}})2\pi e^{1/(6k_{\tilde{B}})})}$
$=\sqrt{(k_{\tilde{B}}+p)/(p^{2}(1-p)2\pi e^{1/(6k_{\tilde{B}})})}$
@@
Now, consider the case $k_{N}=0$. The minimal $B\in\mathbb{N}$ with $\mathfrak{p}^{\mathrm{perm}}_{B}\leq p$ is $B=N-1$. With Proposition 4.1 and $(1-1/N)^{N-1}\geq 1/e$, it follows
@@
since $\sqrt{2\pi e^{1/6}}\geq e$.
Proposition 4.4 shows that for the defined $p$ the only possibility such that $\mathfrak{p}^{\mathrm{perm}}_{B}$ rejects the hypothesis and the binomial strategy does not reject earlier is that $\alpha/\sqrt{2\pi e^{1/6}}<\mathfrak{p}^{\mathrm{perm}}_{B}\leq\alpha$. Therefore, the factor $1/\sqrt{2\pi e^{1/6}}$ can be interpreted as loss we need to incorporate due to the sequential nature of our strategy. Note that some kind of loss is needed, as $\mathfrak{p}^{\mathrm{perm}}_{B}$ is not valid for random stopping times; without the loss factor, Proposition 4.4 would imply that our strategy always rejects earlier than the non-valid $\mathfrak{p}^{\mathrm{perm}}_{B}$. Since $\mathfrak{p}^{\mathrm{perm}}|\{\mathfrak{p}^{\mathrm{lim}}=p^{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}p^{\mathrm{lim}}$, Proposition 4.4 immediately implies the following result.
@@
Let $p=1/\lceil{\sqrt{2\pi e^{1/6}}/\alpha}\rceil$ and $\tau:=\inf\{t\geq 1:W_{t}^{p}\geq 1/\alpha\}$. If $\mathfrak{p}^{\mathrm{lim}}\in[0,p)$, then $\tau$ is almost surely finite and the test $\phi=1\{W_{\tau}^{p}\geq 1/\alpha\}$ has zero resampling risk ($RR_{p^{\mathrm{lim}}}(\phi)=0$).
In Section 5 we propose a strategy that leads to zero resampling risk for all $p^{\mathrm{lim}}\in[0,c)$, where $c$ can be any predefined $c\in(0,\alpha)$.
@@ 4.3 Stopping when the wealth is low
Currently, we assumed that as long as we do not reject the null hypothesis, we will keep generating additional permutations. However, we can also save a lot of permutations when we stop because it is very unlikely that we will reject the null hypothesis if we keep drawing more permutations. More explicitly, we suggest to stop when the current wealth drops below $\alpha$. In this case we would have a test martingale with value less than $\alpha$ and hence the conditional probability of then going on to reject the null hypothesis (crossing wealth $1/\alpha$) would be less than $\alpha^{2}$ under the null. Of course, the conditional probability can be much larger under the alternative, but in this case it is also much more unlikely that we end up with a wealth of $<\alpha$, so it is unlikely that this “stopping out of futility” will hurt power much — we confirm this in simulations. Also note that we can slightly improve our procedure due to this stop. At each step $t$ we know our current wealth $W_{t}$. Therefore, we know whether or not we would stop if we lose in the next round — and when this happens, we might as well bet everything on a win. For the binomial strategy this means that if $W_{t-1}p(t+1)/\mathrm{rank}(Y_{0};Y_{0}^{t-1})<\alpha$, we choose $p=0$ for step $t$. Our final algorithm with the parameter $p$ as in Proposition 4.4 works as follows.
@@
5:     if $W_{t-1}^{p}p(t+1)/\mathrm{rank}_{t-1}<\alpha$ then
@@
11:         $W_{t}^{p}=W_{t-1}^{p}p_{t}(t+1)/\mathrm{rank}_{t-1}$
12:         $\mathrm{rank}_{t}=\mathrm{rank}_{t-1}+1$
@@
14:         $W_{t}^{p}=W_{t-1}^{p}(1-p_{t})(t+1)/(t-\mathrm{rank}_{t-1}+1)$
15:         $\mathrm{rank}_{t}=\mathrm{rank}_{t-1}$
@@ 5 The binomial mixture strategy
As described in Section 1.1, the wealth process obtained by a betting strategy is a test martingale. The evidence contained in multiple dependent test martingales can be combined by averaging them (since averaging preserves the test martingale property). In our case, this means we can average the wealth obtained by various binomial strategies with different parameters $p$. In this section, we show that this can be used to resolve some uncertainty about the parameter choice.
@@ 5.1 Connection to log-optimality
$\bar{W}_{B}^{f}=\int_{0}^{1}W_{B}^{p}f(p)\mu(dp),$
where $W_{B}^{p}$ is the wealth of the binomial strategy after $B$ permutations. By Tonelli’s Theorem, it can be easily verified that $\bar{W}_{t}^{f}$ indeed defines a test martingale with respect to the filtration $(\mathcal{R}_{t})_{t\geq 1}$ (informally, in words: averages of martingales are also martingales). It should be noted that while the individual betting strategies can (and will) depend on the past, the choice of strategies over which we average, in our case given by the density $f$, needs to remain the same over the entire testing process in order to ensure the test martingale property.
@@
The binomial mixture strategy is the log-optimal strategy based on prior $f$ for $\mathfrak{p}^{\mathrm{lim}}$.
@@
Let $W_{B}^{*,f}(k)$ be the wealth of the log-optimal strategy based on prior $f$ after $B$ permutations and $k$ losses and $K_{B}$ denote the random number of losses after $B$ permutations. We show that $W_{B}^{*,f}=\bar{W}_{B}^{f}$ by induction in $B$. For the initial step ($B=1$), simply note that $W_{1}^{*,f}(1)=2\mathbb{E}_{f}[\mathfrak{p}^{\mathrm{lim}}]$ and $W_{1}^{*,f}(0)=2(1-\mathbb{E}_{f}[\mathfrak{p}^{\mathrm{lim}}])$, while $\bar{W}_{1}^{f}(1)=\int_{0}^{1}2pf(p)\mu(dp)=2\mathbb{E}_{f}[\mathfrak{p}^{\mathrm{lim}}]$ and $\bar{W}_{1}^{f}(0)=\int_{0}^{1}2(1-p)f(p)\mu(dp)=2(1-\mathbb{E}_{f}[\mathfrak{p}^{\mathrm{lim}}])$. Next, assume the induction hypothesis (IH) that for some fixed but arbitrary $B\in\mathbb{N}$, it holds that $W_{B-1}^{*,f}(k)=\bar{W}_{B-1}^{f}(k)$ for all $k\in\{0,\ldots,B-1\}$. We will now prove the induction step ($B-1\to B$). First, we consider the case $Y_{B}\geq Y_{0}$. For $k\in\{1,\ldots,B\}$ and $W_{B-1}^{*,f}(k-1)=\bar{W}_{B-1}^{f}(k-1)\neq 0$, it holds
$W_{B}^{*,f}(k)$ $=W_{B-1}^{*,f}(k-1)\frac{B+1}{k}\mathbb{E}_{f}[\mathfrak{p}^{\mathrm{lim}}|K_{B-1}=k-1]$
$=W_{B-1}^{*,f}(k-1)\frac{B+1}{k}\int_{0}^{1}pf_{\mathfrak{p}^{\mathrm{lim}}|K_{B-1}}(p|k-1)\mu(dp)$
$=W_{B-1}^{*,f}(k-1)\frac{B+1}{k}\frac{\int_{0}^{1}pf_{K_{B-1}|\mathfrak{p}^{\mathrm{lim}}}(k-1|p)f(p)\mu(dp)}{\int_{0}^{1}f_{K_{B-1}|\mathfrak{p}^{\mathrm{lim}}}(k-1|p)f(p)\mu(dp)}$
$=W_{B-1}^{*,f}(k-1)\frac{B+1}{k}\frac{\frac{k}{B(B+1)}\int_{0}^{1}W_{B}^{p}(k)f(p)\mu(dp)}{\frac{1}{B}\int_{0}^{1}W_{B-1}^{p}(k-1)f(p)\mu(dp)}$
=$\stackrel{{\scriptstyle\text{IH}}}{{=}}W_{B-1}^{*,f}(k-1)\frac{\bar{W}_{B}^{f}(k)}{W_{B-1}^{*,f}(k-1)}$
$=\bar{W}_{B}^{f}(k).$
@@
A similar calculation can be done for the case $Y_{B}<Y_{0}$ and $k\in\{0,\ldots,B-1\}$. This completes the proof of the induction step, and thus the overall proof. ∎
@@
Theorem 5.1 shows that the log-optimal strategy for a given prior can be obtained by a weighted average over simple binomial strategies. Note that the latter is often easier to compute, since we already derived a simple closed form for the wealth of binomial strategies in Proposition 4.1. Furthermore, it shows that properties of binomial strategies can be transferred to log-optimal strategies, e.g. that the wealth only depends on the number of losses and not the order the losses occur. On the other hand, Theorem 5.1 also shows that every weighted average of binomial strategies can be obtained by single betting strategy in which the bets are defined as in (7) but $p_{t}^{*}$ is replaced with $\mathbb{E}_{f}[\mathfrak{p}^{\mathrm{lim}}|K_{t-1}]$.
5.2 Uniform prior: closed-form wealth is monotone in losses
In practice, we typically do not know the true underlying distribution and thus it is difficult to determine the “best” prior $f$. However, we may reasonably suspect that if the alternative is true, then $\mathfrak{p}^{\mathrm{lim}}$ should be small. Therefore, if we have no information about the true distribution, a reasonable approach is to assume that $\mathfrak{p}^{\mathrm{lim}}$ is uniformly distributed on $[0,c]$ for some $10<c<$ 1. In the following theorem, we derive a simple closed form for the wealth obtained by the corresponding prior
$u_{c}(p)=1\{p\leq c\}/c.$
In this case, $\mu$ is the Lebesgue measure.

### v5

@@ 4.2 A theoretically grounded choice for binomial parameter $p$
Comparing the binomial strategy with the extreme ones in Section 2, the parameter $p$ can be interpreted as the aggressiveness with which we bet. The lower the $p$, the more aggressive we bet on the alternative. In the following, we introduce a choice of $p$ that adapts to the level $\alpha\in(0,1)$ the considered hypothesis is tested at and which offers a good compromise between aggressiveness and safety.
Proposition 4.3.
Let $N=\lceil{\sqrt{2\pi e^{1/6}}/\alpha}\rceil$ and $p=1/N$. If $Q^{\mathrm{perm}}_{T}\leq p$ for some $T\in\mathbb{N}$, then there exists $\tilde{T}\leq T$ such that $W_{\tilde{T}}^{p}\geq 1/\alpha$.
Before we prove Proposition 4.3, we first we need the following lemma.
@@
Let $p=1/N$ for some $N\in\mathbb{N}$, and let $\ell_{N}$ denote the realized number of losses after $N$ permutations. If $p_{O}\leq p$ for some $T\geq N$ and $\ell_{N}\geq 1$, then there exists $\tilde{T}\leq T$ such that $p=\ell_{\tilde{T}}/\tilde{T}$.
Proof of Lemma 4.4.
Let $\tilde{T}=\min\{t\in\mathbb{N}:\ell_{t}/t\leq p,\ell_{t}\geq 1\}$. Note that $\ell_{\tilde{T}}/(N\ell_{\tilde{T}})=p$. Therefore, $N\ell_{\tilde{T}}\leq\tilde{T}$ (otherwise $\ell_{\tilde{T}}/\tilde{T}>p$) and thus $\ell_{N\ell_{\tilde{T}}}\leq\ell_{\tilde{T}}$. Since $\ell_{N\ell_{\tilde{T}}}/(N\ell_{\tilde{T}})\leq\ell_{\tilde{T}}/(N\ell_{\tilde{T}})=p$ and $\ell_{N\ell_{\tilde{T}}}\geq\ell_{N}\geq 1$, we have $\tilde{T}=N\ell_{\tilde{T}}=\ell_{\tilde{T}}/p$. ∎
Proof of Proposition 4.3.
First, assume that $\ell_{N}\geq 1$. Since $\ell_{T}/T\leq Q^{\mathrm{perm}}_{T}$ for all $T\in\mathbb{N}$, Lemma 4.4 shows that there exists $\tilde{T}\leq T$ such that $p=\ell_{\tilde{T}}/\tilde{T}=1/N$. Using the first bound of Proposition 4.2, we obtain
$W_{\tilde{T}}^{p}(\ell_{\tilde{T}})$ $\geq\tilde{T}\sqrt{(\tilde{T}+1)/(\ell_{\tilde{T}}(\tilde{T}-\ell_{\tilde{T}})2\pi e^{1/(6\ell_{\tilde{T}})})}$
$=\sqrt{(\ell_{\tilde{T}}+p)/(p^{2}(1-p)2\pi e^{1/(6\ell_{\tilde{T}})})}$
@@
Now, consider the case $\ell_{N}=0$. The minimal $T\in\mathbb{N}$ with $Q^{\mathrm{perm}}_{T}\leq p$ is $T=N-1$. With Proposition 4.1 and $(1-1/N)^{N-1}\geq 1/e$, it follows
@@
since $\sqrt{2\pi e^{1/6}}\geq e$. ∎
Proposition 4.3 shows that for the defined $p$ the only possibility such that $Q^{\mathrm{perm}}_{T}$ rejects the hypothesis and the binomial strategy does not reject earlier is that $\alpha/\sqrt{2\pi e^{1/6}}<Q^{\mathrm{perm}}_{T}\leq\alpha$. Therefore, the factor $1/\sqrt{2\pi e^{1/6}}$ can be interpreted as loss we need to incorporate due to the sequential nature of our strategy. Note that some kind of loss is needed, as $Q^{\mathrm{perm}}_{T}$ is not valid for random stopping times; without the loss factor, Proposition 4.3 would imply that our strategy always rejects earlier than the non-valid $Q^{\mathrm{perm}}_{T}$. Since $Q^{\mathrm{perm}}_{t}|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}q_{\mathrm{lim}}$ for $t\to\infty$, Proposition 4.3 immediately implies the following result.
@@
Let $p=1/\lceil{\sqrt{2\pi e^{1/6}}/\alpha}\rceil$ and $\tau:=\inf\{t\geq 1:W_{t}^{p}\geq 1/\alpha\}$. If $Q_{\mathrm{lim}}\in[0,p)$, then $\tau$ is almost surely finite and the test $\phi=\mathbbm{1}\{W_{\tau}^{p}\geq 1/\alpha\}$ has zero resampling risk ($RR_{q_{\mathrm{lim}}}(\phi)=0$).
In Section 5 we propose a strategy that leads to zero resampling risk for all $q_{\mathrm{lim}}\in[0,c)$, where $c$ can be any predefined $c\in(0,\alpha)$.
@@ 4.3 Stopping when the wealth is low
Currently, we assumed that as long as we do not reject the null hypothesis, we will keep generating additional permutations. However, we can also save a lot of permutations when we stop because it is very unlikely that we will reject the null hypothesis if we keep drawing more permutations. More explicitly, we suggest to stop when the current wealth drops below $\alpha$. In this case we would have a test martingale with value less than $\alpha$ and hence the conditional probability of then going on to reject the null hypothesis (crossing wealth $1/\alpha$) would be less than $\alpha^{2}$ under the null. Of course, the conditional probability can be much larger under the alternative, but in this case it is also much more unlikely that we end up with a wealth of $<\alpha$, so it is unlikely that this “stopping out of futility” will hurt power much — we confirm this in simulations. Also note that we can slightly improve our procedure due to this stop. At each step $t$ we know our current wealth $W_{t}$. Therefore, we know whether or not we would stop if we lose in the next round — and when this happens, we might as well bet everything on a win. For the binomial strategy this means that if $W_{t-1}p(t+1)/(L_{t-1}+1)<\alpha$, we choose $p=0$ for step $t$. Our final algorithm with the parameter $p$ as in Proposition 4.3 is described in Algorithm 2.
@@
5:     if $W_{t-1}^{p}p(t+1)/(L_{t-1}+1)<\alpha$ then
@@
11:         $W_{t}^{p}=W_{t-1}^{p}p_{t}(t+1)/(L_{t-1}+1)$
12:         $L_{t}=L_{t-1}+1$
@@
14:         $W_{t}^{p}=W_{t-1}^{p}(1-p_{t})(t+1)/(t-L_{t-1})$
15:         $L_{t}=L_{t-1}$
@@ 5 The binomial mixture strategy
As described in Section 1.1, the wealth process obtained by a betting strategy is a test martingale. The evidence contained in multiple dependent test martingales can be combined by averaging them (since averaging preserves the test martingale property). In our case, this means we can average the wealth obtained by various binomial strategies with different parameters $p$. In this section, we show how this can be used to find closed forms for the wealth of the mimicked log-optimal strategy for more complex working priors than $Q_{\mathrm{lim}}=p$.
@@ 5.1 Connection to log-optimality
$\bar{W}_{T}^{f}=\int_{0}^{1}W_{T}^{p}f(p)\mu(dp),$
where $W_{T}^{p}$ is the wealth of the binomial strategy after $T$ permutations. By Tonelli’s Theorem, it can be easily verified that $\bar{W}_{t}^{f}$ indeed defines a test martingale with respect to the filtration $(\mathcal{I}_{t})_{t\geq 1}$ (informally, in words: averages of martingales are also martingales). It should be noted that while the individual betting strategies can (and will) depend on the past, the choice of strategies over which we average, in our case given by the density $f$, needs to remain the same over the entire testing process in order to ensure the test martingale property.
@@
The binomial mixture strategy with density $f$ is equivalent to the log-optimal strategy where $p_{t}^{*}=\mathbb{E}[Q_{\mathrm{lim}}|L_{t-1}]$ is replaced by $p_{t}=\mathbb{E}_{f}[Q_{\mathrm{lim}}|L_{t-1}]$. Hence, if $f$ is the true density of $Q_{\mathrm{lim}}$, then the binomial mixture strategy is log-optimal.
@@
Let $W_{T}^{*,f}(\ell)$ be the wealth of the log-optimal strategy where $p_{t}^{*}$ is replaced by $p_{t}=\mathbb{E}_{f}[Q_{\mathrm{lim}}|L_{t-1}]$ after $T$ permutations and $\ell$ losses and $L_{T}$ denote the random number of losses after $T$ permutations. We show that $W_{T}^{*,f}=\bar{W}_{T}^{f}$ by induction in $T$. For the initial step ($T=1$), simply note that $W_{1}^{*,f}(1)=2\mathbb{E}_{f}[Q_{\mathrm{lim}}]$ and $W_{1}^{*,f}(0)=2(1-\mathbb{E}_{f}[Q_{\mathrm{lim}}])$, while $\bar{W}_{1}^{f}(1)=\int_{0}^{1}2pf(p)\mu(dp)=2\mathbb{E}_{f}[Q_{\mathrm{lim}}]$ and $\bar{W}_{1}^{f}(0)=\int_{0}^{1}2(1-p)f(p)\mu(dp)=2(1-\mathbb{E}_{f}[Q_{\mathrm{lim}}])$. Next, assume the induction hypothesis (IH) that for some fixed but arbitrary $T\in\mathbb{N}$, it holds that $W_{T-1}^{*,f}(\ell)=\bar{W}_{T-1}^{f}(\ell)$ for all $\ell\in\{0,\ldots,T-1\}$. We will now prove the induction step ($T-1\to T$). First, we consider the case $I_{T}=1$. For $\ell\in\{1,\ldots,T\}$ and $W_{T-1}^{*,f}(\ell-1)=\bar{W}_{T-1}^{f}(\ell-1)\neq 0$, it holds
$W_{T}^{*,f}(\ell)$ $=W_{T-1}^{*,f}(\ell-1)\frac{T+1}{\ell}\mathbb{E}_{f}[Q_{\mathrm{lim}}|L_{T-1}=\ell-1]$
$=W_{T-1}^{*,f}(\ell-1)\frac{T+1}{\ell}\int_{0}^{1}pf_{Q_{\mathrm{lim}}|L_{T-1}}(p|\ell-1)\mu(dp)$
$=W_{T-1}^{*,f}(\ell-1)\frac{T+1}{\ell}\frac{\int_{0}^{1}pf_{L_{T-1}|Q_{\mathrm{lim}}}(\ell-1|p)f(p)\mu(dp)}{\int_{0}^{1}f_{L_{T-1}|Q_{\mathrm{lim}}}(\ell-1|p)f(p)\mu(dp)}$
$=W_{T-1}^{*,f}(\ell-1)\frac{T+1}{\ell}\frac{\frac{\ell}{T(T+1)}\int_{0}^{1}W_{T}^{p}(\ell)f(p)\mu(dp)}{\frac{1}{T}\int_{0}^{1}W_{T-1}^{p}(\ell-1)f(p)\mu(dp)}$
=$\stackrel{{\scriptstyle\text{IH}}}{{=}}W_{T-1}^{*,f}(\ell-1)\frac{\bar{W}_{T}^{f}(\ell)}{W_{T-1}^{*,f}(\ell-1)}$
$=\bar{W}_{T}^{f}(\ell).$
@@
A similar calculation can be done for the case $I_{T}=0$ and $\ell\in\{0,\ldots,T-1\}$. This completes the proof of the induction step, and thus the overall proof. ∎
@@
Theorem 5.1 shows that the mimicked log-optimal strategy for a given working prior can be obtained by a weighted average over simple binomial strategies. Note that the latter is often easier to compute, since we already derived a simple closed form for the wealth of binomial strategies in Proposition 4.1. Furthermore, it shows that properties of binomial strategies can be transferred to log-optimal strategies, e.g. that the wealth only depends on the number of losses and not the order the losses occur. On the other hand, Theorem 5.1 also shows that every weighted average of binomial strategies can be obtained by a single betting strategy in which the bets are defined as in (10) but $p_{t}^{*}$ is replaced with $\mathbb{E}_{f}[Q_{\mathrm{lim}}|L_{t-1}]$. Since the binomial mixture strategy with density $f$ is the mimicked log-optimal strategy based on working prior $f$ for $Q_{\mathrm{lim}}$, we also refer to the density $f$ for the binomial mixture strategy as a “working prior” in the following.
5.2 Uniform working prior: closed-form wealth
In practice, we typically do not know the true underlying distribution and thus it is difficult to determine the log-optimal working prior $f$. However, we might know for which values of $Q_{\mathrm{lim}}$ we want our wealth to be large. Therefore, it makes sense to choose our working prior $f$ for $Q_{\mathrm{lim}}$ accordingly, meaning to put most probability mass to these values. For example, if we follow the objective of minimizing the resampling risk (see Section 2.2), then we want a large wealth if $Q_{\mathrm{lim}}\leq\alpha$. Thus, a reasonable approach in this case is to pick $f$ to be uniform on $[0,\alpha]$. In the following theorem, we derive a simple closed form for the wealth obtained by the working prior
$u_{c}(p)=\mathbbm{1}\{p\leq c\}/c,$
where $10<c<$ 1. In this case, $\mu$ is the Lebesgue measure.